# A note on generalized tensor CUR approximation for tensor pairs and tensor triplets based on the tubal product

### v3

@@ 4 Tensor Discrete Empirical Interpolation Method (TDEIM) for lateral/horizontal slice sampling
Theorem 1 shows that the TCUR approximation with the middle tensor defined above can provide approximation within a factor $\tilde{\eta}_{p}+\tilde{\eta}_{q}$ of the best tubal rank $k$ approximation. It also indicates that the conditioning of the problem depends on these two quantities and the lateral/horizontal slices should be selected in such a way that these quantities should be controlled.
@@ 5 Tensor CUR (TCUR) approximation for tensor pairs based on the t-product
Inspired by the success of the MCUR approximation, several generalizations of these approaches to tensors have been proposed. For example, see [34] for the Tucker model, see [35, 36] for the TT decomposition, see [18] for the t-SVD, and see [37] for the TC/TR decomposition. In this note, we focus on the t-SVD case and extend it to a pair of tensors. The intuition behind this generalization is similar to the matrix case. More precisely, the indices of selected lateral slices are the same for two data tensors (see Figure 1 for a graphical illustration). The formal definition of the GTCUR for tensor pairs is presented below.
@@
where $\underline{\bf S}_{1}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\underline{\bf S}_{2}\in\mathbb{R}^{I_{4}\times R\times I_{3}}$ and $\underline{\bf P}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ are index the selection tensors ($R<I_{2}$).
Assume that ${\bf s}_{1},$ and ${\bf s}_{2}$ are the horizontal slice indices of the tensors $\underline{\bf X}$ and $\underline{\bf Y}$, respectively. Also, let ${\bf p}$ be the common lateral slice indices for the both tensors $\underline{\bf X}$ and $\underline{\bf Y}$. Then, formulations (10)-(11) can be equavalently written as
@@
The GTCUR for tensor pairs indeed computes the TCUR approximations of two given data tensors $\underline{\bf X}$ and $\underline{\bf Y}$ and is motivated by the generalized t-SVD (GTSVD). For given data tensors $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\bf Y}\in\mathbb{R}^{I_{4}\times I_{2}\times I_{3}}$, the GTSVD guarantees the following decompositions [17]
@@
where $\underline{\bf U}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}},\,\underline{\bf V}\in\mathbb{R}^{I_{4}\times I_{4}\times I_{3}},\,\underline{\bf C}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}},\,\underline{\bf S}\in\mathbb{R}^{I_{4}\times I_{2}\times I_{3}},\,\underline{\bf Z}\in\mathbb{R}^{I_{2}\times I_{2}\times I_{3}}$. Note that the tensors $\underline{\bf C}$ and $\underline{\bf S}$ are f-diagonal, and the tensors $\underline{\bf U}$ and $\underline{\bf V}$ are orthogonal, while the tensor $\underline{\bf Z}$ is nonsingular. The procedure of the computation of the GTSVD is presented in Algorithm 5. We need to apply the classical GSVD (lines 3-5) to the first $\lceil\frac{I_{3}+1}{2}\rceil$ frontal slices of the tensors $\underline{\bf X}$ and $\underline{\bf Y}$ in the Fourier domain and the rest of the slices are computed easily (Lines 6-12).
@@ 2 $\widehat{\underline{\mathbf{Y}}}={\rm fft}\left({\underline{\mathbf{Y}},[],3}\right)$;
4       $[\widehat{\underline{\mathbf{U}}}_{i},\,\widehat{\underline{\mathbf{V}}}_{i},\,\widehat{\underline{\mathbf{Z}}}_{i},\,\widehat{\underline{\mathbf{C}}}_{i},\,\widehat{\underline{\mathbf{S}}}_{i}]={\rm GSVD}\,(\widehat{\underline{{\bf X}}}(:,:,i),\widehat{\underline{{\bf Y}}}(:,:,i))$;
@@ 12       $\widehat{\underline{\mathbf{S}}}_{i}=\widehat{\underline{\mathbf{S}}}_{I_{3}-i+2}$;
We see that the GTSVD provides the same right tensor $\underline{\bf Z}$ in (14)-(15) and we can use it to sample lateral slices of the data tensors $\underline{\bf X}$ and $\underline{\bf Y}$ based on the TDEIM algorithm. As a result, the same indices can be used to sample the lateral slices for the data tensors $\underline{\bf X}$ and $\underline{\bf Y}$. The horizontal slices of the data tensors $\underline{\bf X}$ and $\underline{\bf Y}$ can also be sampled using the left tensor parts $\underline{\bf U}$ and $\underline{\bf V}$, although they do not necessarily provide identical horizontal slice indices. Following this idea, we can compute the GTSVD of the tensors $\underline{\bf X},\,\underline{\bf Y}$ and by applying the TDEIM to the shared tensor factors $\underline{\bf U},\,\underline{\bf V},\,\underline{\bf Z}$, we can select indices of horizontal and lateral slices of the given data tensors. This approach is summarized in Algorithm 6. In Line 1 of Algorithm 6, we need to compute the GTSVD of two tensors and this is clearly prohibitive for large-scale tensors. However, to tackle this problem, the randomized algorithms proposed in [38] can be used. Note that Lines 10-11 in Algorithm 6 can be efficiently computed, and this is outlined in Algorithm 7.
Input : Two data tensors $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\bf Y}\in\mathbb{R}^{I_{4}\times I_{2}\times I_{3}}$ where $I_{1}\leq I_{2},\,I_{4}\leq I_{2}$ and a target tubal rank $R$
@@
1 $[\underline{\bf U},\underline{\bf V},\underline{\bf Z}]={\rm GTSVD}(\underline{\bf X},\underline{\bf Y})$;   (or randomized GTSVD [38])
@@ 6 ${\underline{\mathbf{G}}}_{4}=\widehat{\underline{\mathbf{Y}}}({\bf s}_{2},:,:)$;
In what follows and similar to [2], we show the connection between the GTCUR of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ and TCUR of the tensors $\underline{\bf X}*\underline{\bf Y}^{-1}$ and $\underline{\bf X}*\underline{\bf Y}^{{\dagger}}$ for non-invertible tensor, or of full rank $\underline{\bf Y}$ in Theorem 3.
@@
    a) If $\underline{\bf Y}$ is invertible, then the selected lateral and horizontal slice indices obtained based on the TCUR of the data tensor $\underline{\bf X}*\underline{\bf Y}^{-1}$ are identical to the index vectors ${\bf s}_{1}$ and ${\bf s}_{2}$ computed from the GTCUR of the tensor pair $(\underline{\bf X},\underline{\bf Y})$.
@@
    b) As a special case, if $\underline{\bf Y}=\underline{\bf I}$, then the GTCUR of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ is the same as the TCUR of the data tensor $\underline{\bf X}$.
@@
    c) If $\underline{\bf Y}$ is noninvertible but of full tubal rank, then the selected lateral and horizontal slice indices obtained via the TCUR of the data tensor $\underline{\bf X}*\underline{\bf Y}^{{\dagger}}$ are identical to the index vectors ${\bf s}_{1}$ and ${\bf s}_{2}$ computed via the GTCUR of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$.
@@
so, $\underline{\bf X}*\underline{\bf Y}^{-1}=\underline{\bf U}*(\underline{\bf C}*\underline{\bf S}^{-1})*\underline{\bf V}^{T}$ because $\underline{\bf Y}^{-1}=\underline{\bf Z}^{-T}*\underline{\bf S}^{-1}*\underline{\bf V}^{T}$. This means that the indices of the selected lateral and horizontal slices obtained via the GTCUR of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ are the same as those obtained from the TCUR of the tensor $\underline{\bf X}*\underline{\bf Y}^{-1}$ as both algorithms use the data tensors $\underline{\bf U}$ and $\underline{\bf V}$ to select the horizontal slices and lateral slices, respectively.
@@ 6 Tensor CUR approximation for tensor triplets based on the t-product
Figure 2: Visualization of the generalized tensor CUR (GTCUR) approximation for tensor triples $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$. The indices of the selected horizontal slices of the tensors $\underline{\bf X}$ and $\underline{\bf Y}$ are identical, but not those for the lateral slices. Also, the indices of the selected lateral slices of $\underline{\bf X}$ and $\underline{\bf Z}$ are the same, but not indices of the selected horizontal slices.
The MCUR approximation for matrix triplets was recently developed in [3] as a generalization of the MCUR. The idea is to use the restricted SVD (RSVD) [7] to select the indices of columns and rows. The RSVD decomposes a matrix relative to two other matrices, and indeed, acts on three matrices simultaneously. This idea can be generalized to tensors based on the t-SVD similarly and we call it GTCUR for tensor triplets. Here, we give the formal definition of this GTCUR approximation.
@@
where ${\bf p},\,{\bf s},\,{\bf s}_{1},\,{\bf p}_{1}$ are the index sets corresponding to the tensor selections $\underline{\bf S}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\widehat{\underline{\bf S}}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\underline{\bf P}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ and $\widehat{\underline{\bf P}}\in\mathbb{R}^{I_{4}\times R\times I_{3}}$.
The intuition behind the GTCUR for tensor triplets also comes from the tensor-based restricted t-SVD (t-RSVD) which is a direct generalization of the classical RSVD [7] to tensors and can be stated as follows. Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, $\underline{\bf Y}\in\mathbb{R}^{I_{1}\times I_{4}\times I_{3}}$, and $\underline{\bf Z}\in\mathbb{R}^{I_{5}\times I_{2}\times I_{3}}$ be given. Generalizing the RSVD to tensors based on the t-product guarantees the following decompositions
@@
where $\underline{\bf U}\in\mathbb{R}^{I_{2}\times I_{2}\times I_{3}},\,\underline{\bf V}\in\mathbb{R}^{I_{5}\times I_{5}\times I_{3}}$ are orthogonal and $\underline{\bf L}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}},\,\underline{\bf W}\in\mathbb{R}^{I_{2}\times I_{2}\times I_{3}}$ are nonsingular. Also, $\underline{\bf D}_{1}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}},\,\underline{\bf D}_{2}\in\mathbb{R}^{I_{1}\times I_{4}\times I_{3}},\,\underline{\bf D}_{3}\in\mathbb{R}^{I_{5}\times I_{2}\times I_{3}}$ have quasi-diagonal frontal slices, which means that if we remove their zero columns and zero rows, they become diagonal. The tensor RSVD (t-RSVD) can be computed through the RSVD of the frontal slices of the underlying data tensors in the Fourier domain. To do so, we can replace the GSVD in Line 4 of Algorithm 5 with the RSVD. It is known that the computation of the RSVD of the frontal slices (matrices) can be performed using a double GSVD as described in [7, 39].
The procedure of the computation of the GTCUR for tensor triples is summarized in Algorithm 8. Lines 6-8 can be efficiently computed in the Fourier domain and similar algorithms like Algorithm 6 can be developed for this computation. The t-RSVD of the tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ provides the common tensor factors $\underline{\bf L},\,\underline{\bf W}$ and extra tensor factors $\underline{\bf U}$ and $\underline{\bf V}$. These bases can be used to sample lateral and horizontal slices of the tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$. Since we have the common tensor factors $\underline{\bf L}$ and $\underline{\bf W}$, the indices, which are used to sample the horizontal slices of the data tensors $\underline{\bf X}$ and $\underline{\bf Z}$ are the same, while the indices for the selection of lateral slices of the data tensors $\underline{\bf X}$ and $\underline{\bf Y}$ are identical. Nevertheless, the tensor bases $\underline{\bf U},\,\underline{\bf V}$ are used to sample lateral slice indices of the tensor $\underline{\bf Y}$ and horizontal slice indices of the tensor $\underline{\bf Z}$. The visualization of this approach is described in Figure 2. So, the idea is to compute the t-RSVD of the tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ and to find the corresponding tensor factors $\underline{\bf U},\,\underline{\bf V},\,\underline{\bf Z},\,\underline{\bf W}$ and use them to find the indices for selecting horizontal and lateral slices using the TDEIM algorithm. We should remark that the computation of the t-RSVD is demanding for the case of large-scale data tensors and the fast randomized GSVD algorithm proposed in [40] can be used for the computation of the double GSVD, which are required to compute the t-RSVD of the tensor triplets. The connection between the GTCUR for tensor triplets with the GTCUR for tensor pairs and the TCUR is described in the next theorem.
@@
    a) If $\underline{\bf Y}$ and $\underline{\bf Z}$ are nonsingular tensors, then the selected lateral and horizontal slice indices from the TCUR decomposition of $\underline{\bf Y}^{-1}*\underline{\bf X}*\underline{\bf Z}^{-1}$ are identical to the index vectors ${\bf p}_{1}$ and ${\bf s}_{1}$, respectively, obtained from a GTCUR decomposition of the tensor triples $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$.
@@
    b) In the particular case where $\underline{\bf Y}=\underline{{\bf I}}$ and $\underline{\bf Z}=\underline{{\bf I}}$, the GTCUR decomposition of $(\underline{\bf X},\underline{\bf I},\underline{\bf I})$ coincides with the TCUR decomposition of $\underline{\bf X}$.
@@
    c) For a special choice of $\underline{\bf Y}=\underline{\bf I}$, a GTCUR approximation of $(\underline{\bf X},\underline{\bf I},\underline{\bf Z})$ coincides with the GTCUR decomposition of $(\underline{\bf X},\underline{\bf Z})$. Similar results can be stated for $\underline{\bf Z}=\underline{\bf I}$.
@@
This means that the selected selected lateral and horizontal slice indices of $\underline{\bf Y}^{-1}*\underline{\bf X}*\underline{\bf Z}^{-1}$ using the TCUR decomposition are the same as the index vectors ${\bf p}_{1}$ and ${\bf s}_{1}$, respectively, obtained from a GTCUR approximation of $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z}).$
b) If $\underline{\bf Y}=\underline{{\bf I}}$ and $\underline{\bf Z}=\underline{{\bf I}}$, then $\underline{\bf L}=\underline{\bf U}*{\underline{\bf D}}_{2}^{-1}$ and $\underline{\bf W}^{T}=\underline{\bf D}_{3}^{-1}*\underline{\bf V}^{T}$. Substituting them in the first part of (21), we have $\underline{\bf X}=\underline{\bf U}*(\underline{\bf D}_{2}^{-1}*\underline{\bf D}_{1}*\underline{\bf D}_{3}^{-1})*\underline{\bf V}^{T}$. This shows that both algorithms use the same tensors $\underline{\bf U}$ and $\underline{\bf V}$ to select the horizontal and lateral slices, respectively. This concludes the proof.
c) If $\underline{\bf Y}=\underline{{\bf I}}$, then from the second part of (21), we have $\underline{\bf L}=\underline{\bf U}*\underline{\bf D}_{2}^{-1}$. If we substitute this in the first and second parts of (21), we arrive at
@@
Clearly, (23) and 24 shows that the GTCUR of the tensor pairs $(\underline{\bf X},\underline{\bf Z})$ provides the same indices for horizontal and lateral slice sampling as the GTCUR decomposition of the tensor pairs $(\underline{\bf X},\underline{\bf I},\underline{\bf Z})$. So, the proof is completed. ∎
@@
Output : A rank-$R$ GTCUR approximation of tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ as
@@
1 Compute t-RSVD (either deterministic or randomized version) of $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ to obtain $\underline{\bf W},\,\underline{\bf Z},\underline{\bf U},\underline{\bf V}$;

### v4

@@ 4 Tensor Discrete Empirical Interpolation Method (TDEIM) for lateral/horizontal slice sampling
Theorem 1 shows that the TCUR approximation with the middle tensor defined above can provide approximation within a factor $\tilde{\eta}_{p}+\tilde{\eta}_{q}$ of the best tubal rank $k$ approximation. It also indicates that the conditioning of the problem depends on these two quantities and the lateral/horizontal slices should be selected in such a way that these quantities should be controlled. In [2], it was empirically shown that these quantities are sufficiently small for many datasets.
@@ 4 $\underline{{\bf R}}=\underline{\bf U}(:,j,:)-\underline{\bf U}(:,1:j-1,:)*\underline{{\bf C}}$;
In the next sections, we extend the TCUR to tensor pairs and tensor triples.
@@ 5 Tensor CUR (TCUR) approximation for tensor pairs based on the t-product
Inspired by the success of the MCUR approximation, several generalizations of these approaches to tensors have been proposed. For example, see [30] for the Tucker model, see [8, 29] for the TT decomposition, see [34] for the t-SVD, and see [12] for the TC/TR decomposition. In this note, we focus on the t-SVD case and extend it to a pair of tensors. The intuition behind this generalization is similar to the matrix case. More precisely, the indices of selected lateral slices are the same for two data tensors (see Figure 2 for a graphical illustration). The formal definition of the GTCUR approximation for tensor pairs is presented below.
@@
where $\underline{\bf S}_{1}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\underline{\bf S}_{2}\in\mathbb{R}^{I_{4}\times R\times I_{3}}$ and $\underline{\bf P}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ are the index selection tensors ($R<I_{2}$).
Assume that ${\bf s}_{1},$ and ${\bf s}_{2}$ are the horizontal slice indices sampled from the tensors $\underline{\bf X}$ and $\underline{\bf Y}$, respectively. Also, let $\bf p$} be the common lateral slice indices for both tensors $\underline{\bf X}$ and $\underline{\bf Y}$. Then, formulations (10)-(11) can be equivalently written as
@@
The GTCUR approximation for tensor pairs indeed computes the TCUR approximations of two given data tensors $\underline{\bf X}$ and $\underline{\bf Y}$ and is motivated by the generalized t-SVD (GTSVD). For given data tensors $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\bf Y}\in\mathbb{R}^{I_{4}\times I_{2}\times I_{3}}$, the GTSVD approximation guarantees the following decompositions [40]
@@
where $\underline{\bf U}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}},\,\underline{\bf V}\in\mathbb{R}^{I_{4}\times I_{4}\times I_{3}},\,\underline{\bf C}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}},\,\underline{\bf S}\in\mathbb{R}^{I_{4}\times I_{2}\times I_{3}},\,\underline{\bf Z}\in\mathbb{R}^{I_{2}\times I_{2}\times I_{3}}$. Note that the tensors $\underline{\bf C}$ and $\underline{\bf S}$ are f-diagonal, and the tensors $\underline{\bf U}$ and $\underline{\bf V}$ are orthogonal, while the tensor $\underline{\bf Z}$ is nonsingular. The process of computing GTSVD is presented in Algorithm 5. We need to apply the classical GSVD (lines 3-5) to the first $\lceil\frac{I_{3}+1}{2}\rceil$ frontal slices of the tensors $\underline{\bf X}$ and $\underline{\bf Y}$ in the Fourier domain and the rest of the slices are computed easily (Lines 6-12).
@@ 2 $\widehat{\underline{\mathbf{Y}}}={\rm fft}\left({\underline{\mathbf{Y}},[],3}\right)$;
4 $\widehat{\underline{\mathbf{U}}}_{i},\,\widehat{\underline{\mathbf{V}}}_{i},\,\widehat{\underline{\mathbf{Z}}}_{i},\,\widehat{\underline{\mathbf{C}}}_{i},\,\widehat{\underline{\mathbf{S}}}_{i}]={\rm GSVD}\,(\widehat{\underline{{\bf X}}}(:,:,i),\widehat{\underline{{\bf Y}}}(:,:,i))$;
@@ 12 $\widehat{\underline{\mathbf{S}}}_{i}=\widehat{\underline{\mathbf{S}}}_{I_{3}-i+2}$;
We see that the GTSVD provides a common right tensor $\underline{\bf Z}$ in (14)-(15) and we can use it to sample lateral slices of the data tensors $\underline{\bf X}$ and $\underline{\bf Y}$ based on the TDEIM algorithm. As a result, the same indices can be used to sample the lateral slices for the data tensors $\underline{\bf X}$ and $\underline{\bf Y}$. The horizontal slices of the data tensors $\underline{\bf X}$ and $\underline{\bf Y}$ can also be sampled using the left tensor parts $\underline{\bf U}$ and $\underline{\bf V}$, although they do not necessarily provide identical horizontal slice indices. Following this idea, we can compute the GTSVD of the tensors $\underline{\bf X},\,\underline{\bf Y}$ and by applying the TDEIM to the shared tensor factors $\underline{\bf U},\,\underline{\bf V},\,\underline{\bf Z}$, we can choose indices for the horizontal and lateral slices of the specified data tensors. This approach is summarized in Algorithm 6. In Line 1 of Algorithm 6, we need to calculate the GTSVD of two tensors, which can be quite challenging for large-scale tensors. However, to address this issue, the randomized algorithms proposed in [4] can be used. Note that Lines 10-11 in Algorithm 6 can be efficiently computed, and this is outlined in Algorithm 7.
Input : Two data tensors $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\bf Y}\in\mathbb{R}^{I_{4}\times I_{2}\times I_{3}}$ where $I_{1}\geq I_{2},\,I_{4}\geq I_{2}$ and a target tubal rank $R$
@@
1 $\underline{\bf U},\underline{\bf V},\underline{\bf Z}]={\rm GTSVD}(\underline{\bf X},\underline{\bf Y})$;  (or randomized GTSVD [4])
@@ 6 ${\underline{\mathbf{G}}}_{4}=\widehat{\underline{\mathbf{Y}}}({\bf s}_{2},:,:)$;
In the subsequent discussion, and akin to [13], we establish the relationship between the GTCUR approximations of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ and the TCUR approximation of the tensors $\underline{\bf X}*\underline{\bf Y}^{-1}$ and $\underline{\bf X}*\underline{\bf Y}^{{\dagger}}$ in Theorem 3.
@@
    a) If $\underline{\bf Y}$ is invertible, then the selected lateral and horizontal slice indices obtained based on the TCUR approximation of the data tensor $\underline{\bf X}*\underline{\bf Y}^{-1}$ are identical to the index vectors ${\bf s}_{1}$ and ${\bf s}_{2}$ computed from the GTCUR approximation of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$.
@@
    b) As a special case, if $\underline{\bf Y}=\underline{\bf I}$, then the GTCUR approximation of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ is the same as the TCUR approximation of the data tensor $\underline{\bf X}$.
@@
    c) If $\underline{\bf Y}$ is noninvertible but of full tubal rank, then the selected lateral and horizontal slice indices obtained via the TCUR approximation of the data tensor $\underline{\bf X}*\underline{\bf Y}^{{\dagger}}$ are identical to the index vectors ${\bf s}_{1}$ and ${\bf s}_{2}$ computed via the GTCUR approximation of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$.
@@
so, $\underline{\bf X}*\underline{\bf Y}^{-1}=\underline{\bf U}*(\underline{\bf C}*\underline{\bf S}^{-1})*\underline{\bf V}^{T}$ because $\underline{\bf Y}^{-1}=\underline{\bf Z}^{-T}*\underline{\bf S}^{-1}*\underline{\bf V}^{T}$. This means that the indices of the selected lateral and horizontal slices obtained via the GTCUR approximation of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ are the same as those obtained from the TCUR approximation of the tensor $\underline{\bf X}*\underline{\bf Y}^{-1}$ as both algorithms use the data tensors $\underline{\bf U}$ and $\underline{\bf V}$ to select the horizontal slices and lateral slices, respectively.
@@ 6 Tensor CUR approximation for tensor triplets based on the t-product
Figure 3: Visualization of the generalized tensor CUR (GTCUR) approximation for tensor triples $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$. The indices of the selected horizontal slices of the tensors $\underline{\bf X}$ and $\underline{\bf Y}$ are identical, but not those for the lateral slices. Also, the indices of the selected lateral slices of $\underline{\bf X}$ and $\underline{\bf Z}$ are the same, but not those for the horizontal slices.
The MCUR approximation for matrix triplets was recently introduced in [15] as an extension of the MCUR concept. The approach involves utilizing the restricted SVD (RSVD) [38] to choose the indices of both columns and rows. The RSVD decomposes a matrix in relation to two other matrices, effectively operating on three matrices at once. This concept can be similarly extended to tensors using the t-SVD, which we refer to as the GTCUR approximation for tensor triplets. Below, we provide the formal definition of this GTCUR approximation.
@@
where ${\bf p},\,{\bf s},\,{\bf s}_{1},\,{\bf p}_{1}$ are the index sets corresponding to the selection tensors $\underline{\bf S}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\widehat{\underline{\bf S}}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\underline{\bf P}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ and $\widehat{\underline{\bf P}}\in\mathbb{R}^{I_{4}\times R\times I_{3}}$.
The intuition behind the GTCUR approximation for tensor triplets also comes from the tensor-based restricted t-SVD (t-RSVD) which is a direct generalization of the classical RSVD [38] to tensors and can be stated as follows. Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, $\underline{\bf Y}\in\mathbb{R}^{I_{1}\times I_{4}\times I_{3}}$, and $\underline{\bf Z}\in\mathbb{R}^{I_{5}\times I_{2}\times I_{3}}$ be given. Generalizing the RSVD to tensors based on the t-product guarantees the following decompositions
@@
where $\underline{\bf U}\in\mathbb{R}^{I_{2}\times I_{2}\times I_{3}},\,\underline{\bf V}\in\mathbb{R}^{I_{5}\times I_{5}\times I_{3}}$ are orthogonal and $\underline{\bf L}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}},\,\underline{\bf W}\in\mathbb{R}^{I_{2}\times I_{2}\times I_{3}}$ are nonsingular. Also, $\underline{\bf D}_{1}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}},\,\underline{\bf D}_{2}\in\mathbb{R}^{I_{1}\times I_{4}\times I_{3}},\,\underline{\bf D}_{3}\in\mathbb{R}^{I_{5}\times I_{2}\times I_{3}}$ have quasi-diagonal frontal slices, which means that if we remove their zero columns and zero rows, they become diagonal. The computation of the RSVD of the matrices can be performed using a double GSVD as described in [38, 9]. Considering this fact, the tensor RSVD (t-RSVD) can be computed through the RSVD of the frontal slices of the underlying data tensors in the Fourier domain. We can indeed modify Algorithm 5 to be used for computing t-RSVD. More precisely, the fft of all three input tensors are computed, then the RSVD is applied for the corresponding frontal slices of the tensors in the Fourier domain. Finally, the inverse fft is applied to get the factor tensors.
[…]
@@
    a) If $\underline{\bf Y}$ and $\underline{\bf Z}$ are nonsingular tensors, then the selected lateral and horizontal slice indices from the TCUR approximation of $\underline{\bf Y}^{-1}*\underline{\bf X}*\underline{\bf Z}^{-1}$ are identical to the index vectors ${\bf p}_{1}$ and ${\bf s}_{1}$, respectively, obtained from a GTCUR approximation of the tensor triples $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$.
@@
    b) In the particular case where $\underline{\bf Y}=\underline{{\bf I}}$ and $\underline{\bf Z}=\underline{{\bf I}}$, the GTCUR approximation of $(\underline{\bf X},\underline{\bf I},\underline{\bf I})$ coincides with the TCUR approximation of $\underline{\bf X}$.
@@
    c) For a special choice of $\underline{\bf Y}=\underline{\bf I}$, a GTCUR approximation of $(\underline{\bf X},\underline{\bf I},\underline{\bf Z})$ coincides with the GTCUR approximation of $(\underline{\bf X},\underline{\bf Z})$. Similar results can be stated for $\underline{\bf Z}=\underline{\bf I}$.
@@
This means that the selected lateral and horizontal slice indices of $\underline{\bf Y}^{-1}*\underline{\bf X}*\underline{\bf Z}^{-1}$ using the TCUR approximation are the same as the index vectors ${\bf p}_{1}$ and ${\bf s}_{1}$, respectively, obtained from a GTCUR approximation of $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z}).$
b) If $\underline{\bf Y}=\underline{{\bf I}}$ and $\underline{\bf Z}=\underline{{\bf I}}$, then $\underline{\bf L}=\underline{\bf U}*{\underline{\bf D}}_{2}^{-1}$ and $\underline{\bf W}^{T}=\underline{\bf D}_{3}^{-1}*\underline{\bf V}^{T}$. Substituting them in the first part of (21), we have $\underline{\bf X}=\underline{\bf U}*(\underline{\bf D}_{2}^{-1}*\underline{\bf D}_{1}*\underline{\bf D}_{3}^{-1})*\underline{\bf V}^{T}$. This indicates that both algorithms utilize the same tensors $\underline{\bf U}$ and $\underline{\bf V}$ for selecting the horizontal and lateral slices, respectively. This completes the proof.
c) If $\underline{\bf Y}=\underline{{\bf I}}$, then from the second part of (21), we have $\underline{\bf L}=\underline{\bf U}*\underline{\bf D}_{2}^{-1}$. If we substitute this in the first part of (21), we arrive at
@@
It is evident that equations (23) and (24) demonstrate that the GTCUR approximation of the tensor pairs $(\underline{\bf X},\underline{\bf Z})$ yields identical indices for both horizontal and lateral slice sampling as the GTCUR approximation of the tensor triples $(\underline{\bf X},\underline{\bf I},\underline{\bf Z})$. This completes the proof.
@@
Output : A tubal rank-$R$ GTCUR approximation of tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ as
@@
1 Compute the t-RSVD (either deterministic or randomized version) of $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ to obtain $\underline{\bf W},\,\underline{\bf Z},\underline{\bf U},\underline{\bf V}$;